OpenStudy (anonymous):

Determine if the sequence 5, 30, 180, 1080... is arithmetic, geometric, or niether.

OpenStudy (anonymous):

If t3-t2=t2-t1 then arithmetic If t3/t2 = t2/t1 then geometrc else neither

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

so its neither?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

then what is it?

OpenStudy (anonymous):

1080/180 =180/30 =30/5

OpenStudy (anonymous):

So, what do u think?

OpenStudy (anonymous):

i dont know?

OpenStudy (anonymous):

geometric?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks ya

OpenStudy (anonymous):

want to help me with more?

OpenStudy (anonymous):

ill be posting more.